# Quantum-inspired activation functions in the convolutional neural network

### v1

@@ Results
Although these quantum circuits have similar expressibility of the Hilbert space and trainability, they have different functional expressions, leading to different performances of the hybrid machine learning approach. Here, we focus on the example of the hybrid quantum-classical CNN. As shown in Fig. 2a, our hybrid CNN includes one convolutional layer with quantum filters, one average pooling layer, one flattened layer, and two fully connected (FC) neural networks (NN). We define the number of features within the first hidden layer as $L$ and set the number of neurons in the last-second hidden layer as $L_{2}=2L$. For the output layer, a SoftMax activation function is employed to facilitate multi-class classification (10 classes are used in our simulations.) alongside the use of a categorical cross-entropy loss function.
The architecture of the quantum filter is delineated in Fig. 2b. For each quantum circuit, we encode an $N_{x}\times N_{y}$ patch $(x_{i,j})$ of the input image into a quantum state by setting the rotation angle in the first $R_{y}$ rotation layer as $\theta_{i\times N_{y}+j}=x_{i,j}\times\pi$. The notations $W$ and $V$ are utilized to denote the controlled-NOT (CNOT) layer within our quantum circuits. When the CNOT layer is omitted between successive $R_{y}$ operators, $V$ is substituted by an identity operator. The angle applied to the subsequent $R_{y}$ operator is a parametric variable, which is used to minimize the loss function. The neuron in the first hidden layer corresponds to the output of the quantum circuit, which is the expectation value of $Z^{\otimes N}$, where $N$ represents the number of qubits, and $Z$ denotes the z-component of the Pauli matrix. Notably, the number of parametric variables within our hybrid CNN aligns with its classical counterpart, except for the biased variable. This configuration facilitates a direct comparison between hybrid and classical CNNs.
We use the MNIST dataset to train our models with $L=16$ and a batch size of 64 [69]. In our work, the training sample size is 57,600, and the testing sample size is 10000. As depicted in Fig. 3, the evolution of the loss function and the accuracy at each epoch during the training phase is plotted for both the five hybrid CNNs and the classical CNN (CCNN). Notably, the analysis reveals that the hybrid CNNs employing quantum circuits 1 and 3 exhibit a markedly accelerated decrease in both the loss function and the residue of the accuracy (defined as 1-accuracy) compared to their counterparts. Hybrid CNNs with QCs 2 and 4 have similar performances and are trained slightly faster than the classical CNN. For the hybrid CNN with QC 5, the loss function and the residue of the accuracy decrease slower than the classical CNN. The better performances of the hybrid CNNs with QCs 1 and 3 show that re-rotating the encoded angle makes better feature selections, which is attributed to the neural network model optimizing the encoding method. In addition, we note that the complex function generated by quantum circuits does not always guarantee a better performance.
The different training speeds among these five hybrid CNNs are attributed to distinct functional expressions of these quantum circuits. To uncover the roots of these variations, we conduct a detailed study of their analytical expressions.
Quantum Circuit 1
@@
We start our analysis with quantum circuit 1. To obtain the output $\langle\Psi|Z^{\otimes N}|\Psi\rangle$ of QC 1, we need to use the following proposition. Here, $|\Psi\rangle$ is the final quantum state of this quantum circuit.
@@
We consider a quantum circuit composed of $N$ qubits, where the initial state $|\psi_{i}\rangle$ of each qubit is described by $|\psi_{i}\rangle=q_{i}|0\rangle+p_{i}|1\rangle$. Upon the application of a CNOT gate between two adjacent qubits first across even-indexed qubits and then odd-indexed qubits, as depicted in Fig. 4a, the expectation value of the operator $Z^{\otimes N}$ is determined by the following recursive relationship:
$\langle\Psi_{1}^{N}|Z^{\otimes N}|\Psi_{1}^{N}\rangle=\left[|p_{1}|^{2}-|q_{1}|^{2}\right]\times\langle\Psi_{3}^{N}|Z^{\otimes N-2}|\Psi_{3}^{N}\rangle,$ (1)
[…]
To demonstrate this proposition, we treat the first two qubits and the remaining qubits, shown in Fig. 4a, separately. The quantum state on the first two qubits after applying a CNOT gate is
@@
[…]
$\otimes|\Psi_{3}^{N}\rangle.$
@@
$f_{QC1}(\theta,\phi)=\left\{\begin{matrix}-\prod^{N}_{i\in\text{odd}}\text{cos}\alpha_{i}&N\in\text{odd}\\ -\text{cos}\alpha_{N}\prod_{i\in\text{odd}}^{N-3}\text{cos}\alpha_{i}&N\in\text{even}\end{matrix},\right.$ (7)
where $\alpha_{i}\!=\!\theta_{i}+\phi_{i}$. Equation (7) shows that the output of QC 1 depends on rotation angles on specific qubits. For example, the output of a $2\times 2$ quantum filter depends on the angle on the first and fourth qubits, which are highlighted by the blue color in Fig. 4b. For a $3\times 3$ filter, the output is determined by the angle on the qubit with odd indexes (see Fig. 4c).
Quantum Circuit 3
Our investigation into the activation function of QC 3 commences by examining the quantum state prior to the final CNOT gate application, denoted as $|\Psi^{N}\rangle$, and illustrated in Fig. 4d. Proposition 2 provides the necessary framework for this analysis, establishing that $|\Psi^{N}\rangle$ can be succinctly described through a recursive formulation, detailed as follows.
@@
[…]
$\left[q_{N}|g^{N-2}\rangle\otimes|0\rangle_{N-1}\!+\!p_{N}|t^{N-2}\rangle\otimes|1\rangle_{N-1}\right]\otimes|1\rangle_{N},$
@@
$|g^{N-2}\rangle\!=\!p_{N-2}|g^{N-4}\rangle\otimes|0\rangle_{N-3}+q_{N-2}|t^{N-4}\rangle\otimes|1\rangle_{N-3},$ (9)
$|t^{N-2}\rangle\!=\!q_{N-2}|g^{N-4}\rangle\otimes|0\rangle_{N-3}+p_{N-2}|t^{N-4}\rangle\otimes|1\rangle_{N-3}.$ (10)
Utilizing the relationship between the output wavefunction $|\Psi\rangle$ of QC 3 and $|\Psi^{N}\rangle$ and Eq. (2), the output $\langle\Psi|Z^{\otimes N}|\Psi\rangle$ of QC 3 can be represented by $|g^{N-2}\rangle$ and $|t^{N-2}\rangle$ via
@@
$\mathrm{cos}\alpha_{N-2}\left[\langle g^{N-4}|Z^{\otimes N-4}|g^{N-4}\rangle+\langle t^{N-4}|Z^{\otimes N-4}|t^{N-4}\rangle\right].$ (13)
Using these recursion equations, the output of QC 3 can be formalized as
@@
$f_{QC3}(\theta,\phi)=\left\{\begin{matrix}\left(\prod_{i=N-2}^{i=i-4,i>4}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{3}\text{cos}\alpha_{1}&N=4+4n\\ \left(\prod_{i=N-2}^{i=i-4,i>6}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{4}\text{cos}\alpha_{2}&N=6+4n\\ \left(\prod_{i=N-2}^{i=i-4,i>5}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{5}\text{cos}\alpha_{2}&N=3+4n\\ \left(\prod_{i=N-2}^{i=i-4,i>5}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{3}\text{cos}\alpha_{1}&N=5+4n,\end{matrix}\right.$

### v2

@@ IV.1 Hybrid quantum-classical convolutional neural network
In the hybrid quantum-classical CNN, the quantum model is adopted in the convolutional operation, as shown in Fig. 2 (a), which includes one convolutional layer with quantum filters, one average pooling layer, one flattened layer, and two fully connected (FC) neural networks (NN). For the output layer, a SoftMax activation function is employed to facilitate multi-class classification alongside the use of a categorical cross-entropy loss function.
In the classical CNN, the convolutional calculation is performed with Tanh or Relu activation functions. In the hybrid CNN, a polynomial function generated from the quantum model replaces this classical activation function. Considering the simplicity of the classical activation function, we assume that a simple function generated from a quantum model can make the hybrid CNN work. The simplest function of polynomial functions is the monotonic function; therefore, we adopt QC 1 in our hybrid CNN. For a nine-qubit circuit, there is no significant difference between $O_{QC1}$ and $O_{QC2}$ (see details in Appendix). Compared to the Tanh function used in the classical CNN, $O_{QC1}$ and $O_{QC2}$ have higher-order nonlinearity and site dependence. To understand the effect of these differences on the performance of the CNN, we perform simulations using five activation functions, as shown in Fig. 2(k). These activation functions include the standard Tanh function (labeled as AF 1), a site-dependent Tanh function (labeled as AF 2), $O_{QC1}$ (labeled as AF 3), a continued product of the cosine series with even indexes (labeled as AF 4), and a continued product of the cosine series without the site dependence (labeled as AF 5).
Our CNNs are adopted to classify images from three datasets: MNIST, Fashion MMNIST (FMNIST), and Letter datasets. For each dataset, 52000 samples with ten classes are used to train the CNNs, and 8000 samples are used for testing. For the CNNs, the number of channels within the first hidden layer is 16, and the number of neurons in the last-second hidden layer is 64. The kernel size is set as $3\times 3$. Figures 2(b)- 2(d) plot the loss function of these five CNNs for three datasets. Compared to the loss function of CNNs using AF 1 and AF 2, the loss functions of CNNs with AF 3 and AF 5 can efficiently converge to a smaller value, implying that the quantum activation functions have higher learnability than the classical activation functions. However, this advantage is not observed in the CNN using AF 4 for the FMNIST dataset, implying that AF 2 is not an optimal activation function for the FMNIST dataset.
We further evaluate the performance of CNNs using AF 1 and AF 3 by analyzing the loss functions and accuracy metrics for both the training and testing datasets, as shown in Figs. 2(e)– 2(j). Although the quantum activation functions exhibit higher learnability, the testing accuracy of CNNs with AF 3 is comparable to that of the standard classical CNNs. We define the optimal model by selecting the epoch corresponding to the minimum loss value on the testing dataset. The table at the bottom of Fig. 2 summarizes the training and testing accuracies, along with the optimal epoch for each of these five CNNs. Compared to the standard classical CNNs, the optimal hybrid CNN models can be obtained with much fewer epochs while achieving similar testing accuracies. These results imply that adopting the quantum activation function can make machine learning more efficient.
It was said that quantum machine learning selects features in the high-dimensional Hilbert space, which has made it shrouded in mystery. However, we note that feature constructions are made by measurements; therefore, understanding feature selections based on measurements is more straightforward than analyzing quantum states. Here, we clarify feature selections using quantum activation functions. Figure 3(a) displays features of trained CNNs using AF 1, AF 3, and AF 4 in the first hidden layer for MNIST, FMNIST, and Letter datasets. Compared to the results of the standard classical CNN, CNNs using quantum activation functions exhibit significantly different features. For the MNIST and the Letter datasets, the model with AF 1 captures comprehensive detail, including both edge delineation and internal patterns within the digit imagery. In contrast, the model leveraging the cosine functions (i.e. AF 3 and AF 4) primarily emphasizes edge information. For the FMNIST dataset, the information inside an object of CNNs using AF 3 and AF 4 does not disappear but is less than that of the CNN using AF 1.
To elucidate this different feature selection, we explore the function $f({\bf\alpha})=\prod_{i}^{3}\text{cos}\alpha_{i}$, where $\alpha_{i}=\theta_{i}+\phi_{i}$. In this formula, $\theta_{i}$ denotes the angle corresponding to the encoded data, and $\phi_{i}$ represents a variable. By considering a three-pixel patch, encoding the background pixels of an image as zero yields an output of zero whenever any $\phi_{i}$ equals $\pi/2$ or $-\pi/2$. This zero output is similarly achieved when encoding a patch inside a digital imagery as $\pi$. At a smooth edge, where $\theta_{i}$ lies between 0 and $\pi$, $f({\bf\alpha})$ is nonzero for $\phi=\pm\pi/2$. Therefore, $f({\bf\alpha})$ can distinguish the edge information and drop redundant information by setting $\phi_{i}=\pm\pi/2$. We examined the trained weight $\phi_{i}$ of the CNN with AF 3 for MNIST, FMNIST, and Letter datasets, plotted in Figs. 3(b)- 3(d). Different colors denote different features. It is found that the weight of each feature has an element close to $\pi/2$. This result demonstrates that our specific quantum activation has the ability to identify the edge information of an object. This ability could disappear if a more complicated quantum activation function is used. With this understanding, we can now clarify the existence of nonzero information inside an object for the FMNIST dataset. The reason is that there is no pixel with a value of zero inside an object; therefore, the value of $\mathrm{cos}(\theta_{i}+\pi/2)$ is no longer zero.
IV.2 Quantum-inspired activation functions
Figure 4: (a)-(c) Accuracy of the convolutional neural network (CNN) using activation functions $f_{1}$, $f_{2}$, $f_{3}$, and AF 1 for the MNIST, FMNIST, and Letter datasets, respectively. (d) Table summarizing the training steps of these four CNNs across these three datasets.
Figure 5: (a) Architecture of the quantum Chebyshev-polynomial network (QCPN). (b) Hybrid QCPN architecture for a two-element input. (c) Three-layer neural network (NN) architecture. (d)-(g) Predictions of the hybrid QCPN model and NN model for four special functions: $y=J_{0}(x)$, $y=J_{1}(x)$, $y=P_{5}(x)$, and $y=P_{6}(x)$, respectively. The dashed curves represent the analytical solutions. (h)-(k) Predictions of the hybrid QCPN model for four special functions: $z=J_{0}(x+y)$, $z=J_{1}(x+y)$, $z=P_{5}(xy)$, and $z=P_{6}(xy)$, respectively. The solid curve denotes the analytical result.
As we explained in Eq. (1), the output of a generic quantum circuit is a polynomial function in which each term is a continued product of the cosine function. In practice, it would be difficult to derive a detailed formula for a complicated quantum circuit. However, it is not necessary to derive such a formula once we understand the mechanism of the hybrid CNNs. Implementing different quantum circuits is equivalent to adopting different polynomial functions. With this understanding, we can directly implement polynomial functions in our classical computer rather than adopting a quantum circuit.
We propose three distinct polynomial functions to initiate a new research direction focused on developing efficient polynomial functions for neural networks. These functions are described as follows:
$f_{1}=$ $\prod_{i\in\text{even}}(\text{cos}(\phi_{i})+\text{cos}(\theta_{i}))+\prod_{i\in\text{odd}}(\text{sin}(\phi_{i})+\text{cos}(\theta_{i}))$
$f_{2}=$ $\prod_{i\in\text{even}}(\text{cos}(\phi_{i})+\text{cos}(\theta_{i}))$ (7)
$f_{3}=$ $\prod_{i}\text{cos}(\phi_{i})\text{cos}(\theta_{i}),$ (8)
where $\theta$ denotes the angle corresponding to the encoded data, and $\phi$ denotes the parametric angle. We note that these three polynomial functions are not derived from any quantum circuit.
Figures 4(a)- 4(c) show the accuracy across three datasets: MNIST, FMNIST, and Letter. The orange bars represent the training data, while the green bars correspond to the testing data. The CNNs utilizing the $f_{1}$ and $f_{2}$ functions achieve performance comparable to the CNN with AF 1. In contrast, the accuracy of the CNN using the $f_{3}$ function is slightly lower across all three datasets. Figure 4 (d) summarizes the number of training steps required to achieve the optimal model for each dataset. Notably, CNNs with quantum-inspired activation functions require fewer training steps compared to the classical CNN using AF 1. These findings demonstrate that manually designing polynomial functions can also support the successful training of CNNs and improve their performance.
IV.3 Quantum Chebyshev-polynomial network
Higher-order nonlinear activation functions can not only reduce the number of training steps but also decrease the depth of deep learning models. In this section, we develop a quantum Chebyshev-polynomial network leveraging the properties of quantum activation functions to demonstrate the efficiency of higher-order nonlinear activation functions.
As is well known that any continuous function $F(x)$ defined over the interval $[a,b]$ can be approximated using a Chebyshev polynomial as follows:
$F(x)=\sum_{n=0}^{\infty}c_{n}T_{n}(L(x))+\epsilon,$ (9)
where $L(x)$ is a linear transformation that maps $x$ to the interval $[-1,1]$, $T_{n}(\xi)$ is the Chebyshev polynomial of the first kind, and $\epsilon$ is a small error term. Alternatively, $F(x)$ can can be expanded using a Fourier series. However, the Chebyshev polynomial expansion has the advantage that the coefficients $c_{n}$ rapidly decrease as $n$ increases, allowing us to approximate $F(x)$ accurately with only a finite number of terms.
For a machine learning task where the target value is a continuous function of the input, we can use Chebyshev polynomials to establish a relationship between the target value and the input. Based on this concept, we develop a quantum Chebyshev-polynomial network (QCPN), as illustrated in Fig. 5(a). The QCPN is constructed using parametric $R_{y}$ rotation gates and controlled-$R_{y}(\xi)$ gates, where $\xi=\mathrm{arccos}(x)$. Here, we assume the input $x$ lies within the range of [-1, 1]. The measurement is performed by applying the Pauli $Z$ gate on the first and second qubits, while the identity operator is applied to the remaining qubits. For an $n$-qubit ($n>2$) circuit, the output of the QCPN can be expressed as
$O(x)=\sum_{i}^{(n-1)(n-2)}(-1)^{i}a_{i}^{2}T_{i}(x),$ (10)
where $a_{i}$ is determined by the angle in the parametric $R_{y}$ gates. Since the $a_{i}$ values are not independent of each other, multiple QCPNs can be employed to increase the degrees of freedom. Consequently, the target value $y(x)$ can be expressed as:
$y(x)=\sum_{i}w_{i}O_{i}(x)+b,$ (11)
where $O_{i}$ denotes the output of the $i$-th QCPN, and $w_{i}$ represents the corresponding weight.
When the input consists of $m$ elements, the target value can be approximated as
$y(x_{1},x_{2},\cdots,x_{m})=\sum_{i}w_{i}\prod_{j}^{m}O_{i,j}(x_{j})+b.$ (12)
If $y(x_{1},x_{2},\cdots,x_{m})$ is a separable function, i.e. $y(x_{1},x_{2},\cdots,x_{m})=\prod_{j}^{m}y_{j}(x_{j})$, then Eq. (12) naturally holds. Figure 5(b) illustrates the architecture of the hybrid quantum network for an input with two elements. The notation $A\times B$ in Fig. 5(b) represents the multiplication of the outputs from quantum circuits with inputs $x_{1}$ and $x_{2}$. Equation (12) implies that we can approximate any continuous function using a hybrid QCPN with three layers. In contrast, a standard three-layer neural network cannot achieve this level of expressivity, which will be discussed later.
We first train the hybrid QCPN to predict the zeroth and first orders of the Bessel function, $J_{0}(x)$ and $J_{1}(x)$, as well as the fifth and sixth orders of the Legendre function, $P_{5}(x)$ and $P_{6}(x)$. The first hidden layer of the hybrid QCPN has two neurons generated from two five-qubit quantum circuits. For comparison, we also train a three-layer neural network (NN) to predict these functions. The architecture of the NN is shown in Fig. 5(c), where the first layer employs the Tanh activation function, and the second layer uses a linear function. Both networks are trained using 50,000 samples and tested with 1,000 samples. Figures 5(d)- 5(g) show the predictions from both models. The red symbols represent the prediction of the hybrid QCPN, and the blue symbols indicate the prediction of the NN, with the dashed curve representing the exact values of these special functions. The hybrid QCPN model achieves high prediction accuracy for all four functions. In contrast, the NN model fails to accurately predict $J_{0}$ and $J_{1}$ as $x>5$ and $P_{5}(x)$ and $P_{6}(x)$. Increasing the number of neurons in the second layer of the NN does not improve accuracy, whereas adding one or two more layers is necessary for the NN to succeed. These results demonstrate that the hybrid QCPN, with its higher-order nonlinear activation functions, can reduce the depth of a neural network, thereby decreasing the overall model complexity.
Furthermore, we examine Eq. (12) by training the hybrid QCPN on four additional special functions, $z=J_{0}(x+y)$, $z=J_{1}(x+y)$, $z=P_{5}(xy)$, and $z=P_{6}(xy)$. The curves of these special functions are shown in Figs. 5(h)- 5(k). For each QCPN, we use a five-qubit configuration, while the classical layer consists of ten input nodes. The red, green, and blue circles in Figs. 5(h)- 5(k) represent the predicted values of the trained hybrid QCPN for the testing dataset. The model achieves high accuracy in predicting these special functions, indicating that the hybrid QCPN can effectively tackle complex machine-learning tasks involving input with multiple elements.
Similarly, this method can be adapted to classical neural networks to enhance their performance. A related approach was recently proposed in the context of Chebyshev polynomial-based Kolmogorov-Arnold Networks [66]. Unlike the method in Ref. [66], the QCPN can automatically express a continuous function using a polynomial with multiple terms, eliminating the need to construct Chebyshev bases manually. In addition, the QCPN can build a Chebyshev function up to the $(n-1)(n-2)$-th order using only $n$ qubits, which is more efficient than the classical method. Furthermore, we utilize a different polynomial function to describe continuous functions with high-dimensional inputs.
V Perspective
Our work investigates the functional expressibility of quantum circuits for machine learning tasks, providing insights distinct from those in Ref. [67], which primarily explores the capacity of simulating the quantum state of a quantum circuit. Unlike classical neural networks, quantum neural networks leverage specialized activation functions that can potentially reduce both the number of training steps and the model size. Interestingly, similar benefits can be achieved on classical computers by employing quantum-inspired activation functions. This finding raises an important question: Do we really need to use quantum computers to achieve these benefits?
The answer is yes! Here, we aim to discuss the benefits of leveraging quantum circuits and quantum computers for machine-learning tasks. In our work, we manually design a quantum-inspired activation function to extract features from input images. This construction is relatively straightforward when the quantum circuit is small. However, as the size of the quantum model scales to hundreds or thousands of qubits, constructing an efficient polynomial function for machine learning tasks becomes increasingly challenging. Adopting a quantum circuit can significantly simplify the process of building such polynomial functions. For instance, one can repeatedly apply specific quantum layers within a quantum circuit to create complex quantum activation functions. However, developing such large-scale quantum models depends on advances in quantum computing hardware, which would require highly reliable quantum computers with at least 1,000 logical qubits. Recent developments, such as IBM’s error-correction techniques utilizing fewer physical qubits, have shown promise in realizing quantum computers with a large number of logical qubits [68].
In addition, developing quantum machine learning techniques is crucial for addressing the challenges inherent in quantum systems. For example, a quantum machine learning model can be directly employed to mitigate quantum errors in quantum computers [69]. Moreover, quantum machine learning models can be used to classify the quantum states of a system simulated on quantum computers [70]. These applications highlight the importance of quantum machine learning techniques in solving quantum many-body physics problems.
Understanding the mathematical expression of quantum circuits also motivates us to explore quantum-inspired activation functions. As we discussed, constructing an efficient polynomial function with hundreds or thousands of terms is challenging. It would be interesting to explore some rules to guide us to construct polynomial functions with high learning capability. If such a rule can be derived, developing a quantum model for classical machine learning tasks would become less critical.
VI Conclusion
Our work explored the functional expressibility of a generic quantum circuit. By examining two shallow quantum circuits integrated into a convolutional neural network, we demonstrated that incorporating a quantum activation function can enhance model performance by reducing the required number of training steps. Additionally, we investigated three quantum-inspired activation functions and found that they can achieve prediction accuracies comparable to those of standard classical CNNs. Inspired by the properties of Chebyshev polynomial functions, we developed a hybrid quantum Chebyshev-polynomial network capable of approximating any continuous function using just three layers. In contrast, a standard three-layer neural network fails to predict continuous functions with high-order nonlinearities.
We provided an in-depth discussion on future research directions in quantum machine learning. We advocate for efforts to optimize large-scale quantum circuits and to explore systematic methods for constructing efficient polynomial functions with thousands of terms. Advancements in either of these areas could significantly enhance machine learning model performance by reducing the number of training steps and model complexity. This is particularly crucial for large-scale natural language models like ChatGPT and Gemini.
Acknowledgement
This work is supported by the National Center for Transportation Cybersecurity and Resiliency (TraCR) (a U.S. Department of Transportation National University Transportation Center) headquartered at Clemson University, SC, USA. Any opinions, findings, conclusions, and recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of TraCR. The U.S. Government assumes no liability for the contents or use thereof. We acknowledge the computational support from Palmetto, a supercomputing cluster at Clemson University.
Figure 6: (a) The quantum state $|\Psi_{1}^{N}\rangle$ of quantum circuit 1 is divided into two parts, $|\Psi_{1}^{2}\rangle$ and $|\Psi_{3}^{N}\rangle$, where $N$ is the total number of qubits. (b) The contribution of a $2\times 2$ filter encoded into quantum circuit 1 comes from the first and fourth qubits. (c) The contribution of a $3\times 3$ filter encoded into quantum circuit 1 comes from qubits with odd indexes. (d) The analysis of quantum circuit 3. (e) The contribution of a $2\times 2$ filter encoded into quantum circuit 3 comes from the first and third qubits. (f) The contribution of a $3\times 3$ filter encoded into quantum circuit 3 comes from the first, third, and seventh qubits.
Appendix
The function expression of quantum circuit 1
@@ Acknowledgement
To obtain the output $\langle\Psi|Z^{\otimes N}|\Psi\rangle$ of QC 1, we need to use the following proposition. Here, $|\Psi\rangle$ is the final quantum state of this quantum circuit.
@@
We consider a quantum circuit composed of $N$ qubits, where the initial state $|\psi_{i}\rangle$ of each qubit is described by $|\psi_{i}\rangle=q_{i}|0\rangle+p_{i}|1\rangle$. Upon the application of a CNOT gate between two adjacent qubits first across even-indexed qubits and then odd-indexed qubits, as depicted in Fig. 6(a), the expectation value of the operator $Z^{\otimes N}$ is determined by the following recursive relationship:
$\langle\Psi_{1}^{N}|Z^{\otimes N}|\Psi_{1}^{N}\rangle=\left[|p_{1}|^{2}-|q_{1}|^{2}\right]\times\langle\Psi_{3}^{N}|Z^{\otimes N-2}|\Psi_{3}^{N}\rangle,\$ (13)
where $|\psi_{n}^{m}\rangle=\otimes_{i=n}^{i=i+2,i<m}\text{CX}(i,i+1)\otimes_{j=n+1}^{j=j+2,j<m}\text{CX}(j,j+1)\otimes_{i=n}^{m}|\psi_{i}\rangle$. The notation $\text{CX}(i,j)$ denotes a CNOT gate with qubit $i$ as the controller and qubit $j$ as the target.
To demonstrate this proposition, we treat the first two qubits and the remaining qubits, shown in Fig. 6(a), separately. The quantum state on the first two qubits after applying a CNOT gate is
@@
$|\Psi_{1}^{N}\rangle\!=\!$ $\big{[}q_{1}q_{2}|00\rangle+q_{1}p_{2}|01\rangle X_{3}+p_{1}q_{2}|11\rangle X_{3}+p_{1}p_{2}|10\rangle\big{]}$
$\otimes|\Psi_{3}^{N}\rangle.$ (15)
@@
$O_{QC1}(\theta,\phi)=\left\{\begin{matrix}-\prod^{N}_{i\in\text{odd}}\text{cos}\alpha_{i}&N\in\text{odd}\\ -\text{cos}\alpha_{N}\prod_{i\in\text{odd}}^{N-3}\text{cos}\alpha_{i}&N\in\text{even}\end{matrix},\right.$ (17)
where $\alpha_{i}\!=\!\theta_{i}+\phi_{i}$. Equation (17) shows that the output of QC 1 depends on rotation angles on specific qubits. For example, the output of a $2\times 2$ quantum filter depends on the angle on the first and fourth qubits, which are highlighted by the blue color in Fig. 6(b). For a $3\times 3$ filter, the output is determined by the angle on the qubit with odd indexes (see Fig. 6(c)).
The function expression of quantum circuit 2
Our investigation into the activation function of QC 2 commences by examining the quantum state prior to the final CNOT gate application, denoted as $|\Psi^{N}\rangle$, and illustrated in Fig. 6(d). Proposition 2 provides the necessary framework for this analysis, establishing that $|\Psi^{N}\rangle$ can be succinctly described through a recursive formulation, detailed as follows.
@@
$|\Psi^{N}\rangle=\left[p_{N}|g^{N-2}\rangle\otimes|0\rangle_{N-1}+q_{N}|t^{N-2}\rangle\otimes|1\rangle_{N-1}\right]\otimes$
$|0\rangle_{N}+\left[q_{N}|g^{N-2}\rangle\otimes|0\rangle_{N-1}+p_{N}|t^{N-2}\rangle\otimes|1\rangle_{N-1}\right]\otimes|1\rangle_{N},$ (18)
@@
$|g^{N-2}\rangle=p_{N-2}|g^{N-4}\rangle\otimes|0\rangle_{N-3}+q_{N-2}|t^{N-4}\rangle\otimes|1\rangle_{N-3},$
$|t^{N-2}\rangle=q_{N-2}|g^{N-4}\rangle\otimes|0\rangle_{N-3}+p_{N-2}|t^{N-4}\rangle\otimes|1\rangle_{N-3}.$
Utilizing the relationship between the output wavefunction $|\Psi\rangle$ of QC 2 and $|\Psi^{N}\rangle$ and Eq. (2), the output $\langle\Psi|Z^{\otimes N}|\Psi\rangle$ of QC 2 can be represented by $|g^{N-2}\rangle$ and $|t^{N-2}\rangle$ via
@@
$\mathrm{cos}\alpha_{N-2}\left[\langle g^{N-4}|Z^{\otimes N-4}|g^{N-4}\rangle+\langle t^{N-4}|Z^{\otimes N-4}|t^{N-4}\rangle\right].$
Using these recursion equations, the output of QC 2 can be formalized as
@@
$O_{QC2}(\theta,\phi)=\left\{\begin{matrix}\left(\prod_{i=N-2}^{i=i-4,i>4}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{3}\text{cos}\alpha_{1}&N=4+4m\\ \left(\prod_{i=N-2}^{i=i-4,i>6}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{4}\text{cos}\alpha_{2}&N=6+4m\\ \left(\prod_{i=N-2}^{i=i-4,i>5}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{5}\text{cos}\alpha_{2}&N=3+4m\\ \left(\prod_{i=N-2}^{i=i-4,i>5}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{3}\text{cos}\alpha_{1}&N=5+4m,\end{matrix}\right.$ (24)